1. Use the formula P(A) PABP(B) + P(AlBc)P(B) to prove that if P(AB) P (AlBc) then A and B are independent. Then prove the converse (that if A and B are independent then P(AIB)- P(ABe). [Assume that P(B) > 0 and P(B) > 0.]
Answer #1

P(A)= P(A|B)P(B)+ P(A|B^c)P(B^c)

i) If P(A|B) = P(A|Bc),

Rightarrow P(A)= P(A|B)P(B)+ P(A|B)P(B^c)

Rightarrow P(A)= P(A|B)[P(B)+P(B^c)]

P(A) P(AB)1

P(A) P(AB)

That implies, A and B are independent.

ii)Since A and B are independent, that implies PA)-PAB) and also P(A) = P(A|B^c)

Hence, we can write P(A|B) = P(A|B^c)